A radioactive element sodium-24 has a half-life of 14 hours and is used to help locate obstructions in blood flow. A procedure requires 1.5 grams and is scheduled to take place in 37 hours.
At what rate is the sodium-24 decaying?
What is the minimum amount of sodium-24 that must be on hand now?

Answer by stanbon(75887)

A radioactive element sodium-24 has a half-life of 14 hours and is used to help locate obstructions in blood flow. A procedure requires 1.5 grams and is scheduled to take place in 37 hours.
---------
EQUATION:
A(t) = Ao(1/2)^(t/14)
1.5 = Ao(1/2)^(37/14)
Ao = 1.5 / [(1/2)^(37/14)]
Ao = 1.5 / 0.16011...
Ao = 9.369 grams
--------------------
At what rate is the sodium-24 decaying?
[9.369-1.5]/37 = 0.21266 grams per hour
-----------------------------
What is the minimum amount of sodium-24 that must be on hand now?
9.369 grams
==============
Cheers,
Stan H.